OpenStudy (16shuston):

Is this right? i got the answer of 21.4

OpenStudy (16shuston):

http://prntscr.com/7xktvm

OpenStudy (biohazard9064):

are there answers or just a blank?

OpenStudy (16shuston):

i got 21.4 as an answer

OpenStudy (biohazard9064):

yes but is it multiple choice or just fill in the bank

OpenStudy (biohazard9064):

Well I got d≈27.53in

OpenStudy (16shuston):

its fill in blank

OpenStudy (biohazard9064):

because d=w^2+l^2=22.4^2+16^2≈27.52744in

OpenStudy (biohazard9064):

so it would be d≈27.53in But d≈27.50in because to the nearest tenth

OpenStudy (16shuston):

okay i see now, thanks

OpenStudy (biohazard9064):

np